HanAkoSolo (jamierox4ev3r):

Find the inverse of the function \(\huge\frac{4x-3}{2x+1}\)

HanAkoSolo (jamierox4ev3r):

Once again, I need someone to check out what I've done. So first, I've multiplied both sides by 2x+1 As a result, I got this: x(2y+1)=4y-3

HanAkoSolo (jamierox4ev3r):

Then, I distributed the x into the parentheses 2yx+x=4y-3

HanAkoSolo (jamierox4ev3r):

Then, I subtracted x and 4y from both sides, which left me with this: 2yx-4y=-x-3

HanAkoSolo (jamierox4ev3r):

from there, I condensed 2yx-4y, so the equation was left as follows: 2y(x-2) And the expression as a whole looked like this: 2y(x-2)=-x-3

HanAkoSolo (jamierox4ev3r):

from here, I divided both sides by x-2. The point here is to isolate the y. Problem is, there is a two right in front of the y. Does this mean that I have to divide again by 2 in order to fully isolate the y?

OpenStudy (ankit042):

Looks good to me! you should get (x+3)/2(2-x)

HanAkoSolo (jamierox4ev3r):

Wait. So it would look like \(\huge\frac{x+3}{2(2-x)}\)

OpenStudy (ankit042):

Yeah!

HanAkoSolo (jamierox4ev3r):

I don't quite understand how you would get from this: \(\huge 2y(x-2)=-x-3\) to this: \(\huge\frac{x+3}{2(2-x)}\)

HanAkoSolo (jamierox4ev3r):

The last clear step to me is dividing both sides by x-2 that would give you \(\huge 2y=\frac{-x-3}{x-2}\)

HanAkoSolo (jamierox4ev3r):

\(\huge \color{red}2y=\frac{-x-3}{x-2}\) This 2 is eluding me so much and I have no clue why.

OpenStudy (cggurumanjunath):

multiply - sign on both sides !

OpenStudy (cggurumanjunath):

negative sign

OpenStudy (tkhunny):

Why do you doubt?

HanAkoSolo (jamierox4ev3r):

I'm not sure xD it's just when I see this: \(\huge 2y=\frac{-x-3}{x-2}\) I think that the whole thing is going to end up looking like this \(\huge y=\frac{\frac{-x-3}{x-2}}{2}\)

HanAkoSolo (jamierox4ev3r):

And that kind of seems super messy to me. I have no idea how to make it appear clean and neat like ankit made it seem to be. I'm sure he's right, i just don't know how to get there

jimthompson5910 (jim_thompson5910):

@Jamierox4ev3r dividing both sides by 2 is the same as multiplying both sides by 1/2 \[\Large 2x = 4\] \[\Large {\color{red}{\frac{1}{2}}}*2x = {\color{red}{\frac{1}{2}}}*4\]

OpenStudy (tkhunny):

\(\dfrac{\dfrac{a}{b}}{2} = \dfrac{a}{b}\cdot\dfrac{1}{2} = \dfrac{a}{2b}\) \(-x+2 = -(x-2) = 2-x\) No mysteries in there.

HanAkoSolo (jamierox4ev3r):

Woah I did not see it that way. That makes sense, so basically it's like doing this: \(\huge\frac{-x-3}{x-2}\times\frac{1}{2}\)

HanAkoSolo (jamierox4ev3r):

still don't get how the numerator becomes positive

OpenStudy (ankit042):

Yeah! now multiply both numerator and denominator by -1

OpenStudy (tkhunny):

\(-\dfrac{a}{b} = -\dfrac{-a}{-b} = \dfrac{-a}{b} = \dfrac{a}{-b}\) Look around. It's there.

HanAkoSolo (jamierox4ev3r):

I apologize, but I still don't see why the numerator and denominator should be multiplied by -1. If i multiply -x-3 / x-2 by 1/2, then wouldn't the final answer be \(\huge\frac{-x-3}{2(x-2)}\)

OpenStudy (mathstudent55):